A person is driving a car with speed of vo. The person brakes to stop over time t and distance d. The magnitude of the person acceleration is a. If the person's acceleration were tripled, how long would it take her to stop?
1
Expert's answer
2020-09-21T08:30:37-0400

The time of motion of the to the rest

"t_1=\frac{v_0}{a_1}"

If the person's acceleration was tripled, the time of motion would be

"t_2=\frac{v_0}{a_2}=\frac{v_0}{3a_1}=\frac{1}{3}t_1"